OpenStudy (anonymous):

x + y = 1120 3x - 2y = 1300 ?????????????? Help/:

OpenStudy (natalieee):

what are you asking ?

OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1120. If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, how many two-step equations auditioned today?

OpenStudy (anonymous):

that was question and I typed it out in a two-step equation

OpenStudy (natalieee):

sorry im not good with these.

OpenStudy (anonymous):

use substitution method

OpenStudy (anonymous):

do u knw?

OpenStudy (anonymous):

kind of?

OpenStudy (anonymous):

but I've gotten it wrong/:

OpenStudy (mathstudent55):

Multiply first eq by 2 and add to second eq 2x + 2y = 2240 3x - 2y = 1300 -------------- 5x = 3540 x = 708 x + y = 1120 708 + y = 1120 y = 412